how much work is done in raising a stone of mass 5 kg and relative density 3 lying at the bed at the of a lake through height of 3 meter? (taking g = 10 10m/s square )?????


(a) 25 j              (b) 100j


(c) 75j                (d) none of the above....

Grade:11

1 Answers

vikas askiitian expert
13 years ago

net force = W - boyancy                  

ds , dw are dencity  of stone & water respectively & V  is volume of stone then ,

boyancy = dwVg

W = mg = dsVg

Fnet = Vg(ds-dw)

Fnet = ma = dsVg

a = g(1-dw/ds

relative dencity = ds/dw = 2 (given)

so  ,a = g/2

now work is done against this force

work = maX                 (X displacement)

       =5*(g/2)3 = 75j